# Perturbation-based Effect Measures for Compositional Data

### v3

@@ 4.1.1 Effects of diversity on income in semi-synthetic US census data
We now use this raw individual-level data to construct community-level datasets by artificially grouping individuals together. The grouping is deliberately constructed in a way such that diversity affects income and this relationship is confounded by the remaining variation of the racial composition. This semi-synthetic example is only intended as an illustration. A genuine analysis of this data would require knowledge of the true community structure, which is not available in the data extract we are using here. We emphasize that even if the individual-level data is available, one may still be interested in effects of the racial composition of the community an individual lives in on that individual’s income. For such an analysis, one can add the racial composition of an individual’s community as a variable and then perform the same analysis as here but now on the individual-level data while adjusting for individual-level confounding factors. Importantly, even for such an individual-level analysis, the simplex confounding can bias the results in the same way as we illustrate here. We provide an example of this bias occurring on the individual-level effect in Section S4.5 of the supplementary material.
@@
We group all individual-level observations into groups of around $50$ individuals. We then create a new response variable $Y\in[0,1]$ by averaging the individual-level compensation in each group. Similarly, we create a predictor variable $X\in\mathbb{R}\times[0,1]\times\{1,\dots,16\}$ by averaging age and sex and selecting the majority class of the categorical variable education (breaking ties randomly). Finally, we create a compositional predictor $Z\in\Delta^{2}$ by computing the proportions of ‘White’, ‘Black’ and ‘Other’ individuals in each group. To ensure, a (partially) known effect of increased racial diversity on compensation, we construct three categories of groups (0, 1 and 2). We construct these categories in two different ways:
@@
  • Grouping by race and education: Groups in category 1 are sampled to be similarly diverse and with lower educational levels than the total population and category 2 groups are sampled to be more diverse and with higher educational levels than the total population. The groups in category 0 consist of groups constructed randomly from the remaining individuals.
  • Grouping by race and compensation: Groups in category 1 are sampled to be similarly diverse and with lower average compensation than the total population and category 2 groups are sampled to be more diverse and with higher average compensation than the total population. The groups in category 0 consist of groups constructed randomly from the remaining individuals.
@@
We then sample groups from each category, using a weighted sampling approach described in detail in Section S4.3 of the supplementary material. This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation. For the data where we have grouped on education the racial composition for the resulting groups is illustrated in the left plot of Figure 4. We expect that – by construction – there is a positive effect of increased diversity on compensation because education is positively associated with compensation. This can be seen in the right plot of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive. The same occurs when grouping on compensation and the corresponding figure can be found in Section S4.4 of the supplementary material.
For each type of grouping the racial composition for the resulting groups is illustrated in the top half of Figure 4. In both cases we expect that – by construction – there is a positive effect of increased diversity on compensation (for the grouping by race and education this is because education is positively associated with compensation). This can be seen in the lower half of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
Figure 4: Summary of the semi-synthetic ‘Adult’ dataset when grouping on education. Left: Racial compositions in a ternary plot. Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red). Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle). Depending on the color the dependence between $Y$ and the Gini coefficient changes. The Gini coefficient represents a similarity rather than a diversity measure.
@@
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use minus the Gini coefficient $L=-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 5, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3 where random forests are used for nuisance function estimation when including additional covariates. The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded (e.g., educational levels may well be correlated with race and average compensation in actual communities).
@@
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
@@ 4.1.2 Effects of diversity on grades in New York schools
In this section we analyze data from a 2018 Kaggle competition containing data on schools in New York (PASSNYC, 2018). The data (collected in 2016) contains average grades in English Language Arts (ELA) and in math for 1217 schools along with additional background information on the school such as a summary of the economic need of the students, percentage of students with English as second language, racial composition of the school and various indicators of school performance (for a full list of control variables and our pre-processing of these, see Section S4.6 of the supplementary material). We let $X$ denote all background variables except racial composition which we denote by $Z$. Race is given as whole percentages of ‘Asian’, ‘Black’, ‘Hispanic’ and ‘White’ students which we rescale so that $Z\in\Delta^{3}$. We are interested in the effect of diversity, again measured by the Gini coefficient, on the average grade of the schools (either ELA or math). We repeat the analysis in Section 4.1.1 on this real data and the results can be seen in Table 4.
@@
$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.345}}$ $(0.105,0.585)$ $0.098$ $(-0.175,0.371)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.163}}$ $(0.002,0.323)$ $0.047$ $(-0.167,0.261)$
@@ 4.2 Variable influence measures when predicting BMI from gut microbiome
We now analyze different types of average perturbation effects and compare them with the commonly used log-contrast based analysis. We fit both a standard log-contrast regression and an $\ell^{1}$-penalized log-contrast regression and then use the coefficients as feature influences. Since the log-contrast model is only defined on the open simplex, that is, it does not permit zeros in the data, one usually adds a ‘pseudocount’ to each observation before rescaling to the simplex. We follow a standard recommendation and use the minimum non-zero observation of $Z$ divided by $2$ (this is not required for the perturbation-based measures). To investigate how the log-contrast approach compares to the perturbation-based approach, we plot the most significant hits using $\mathrm{CFI}_{\mathrm{mult}}$ (with random forests as regression method) and a log-contrast regression model in the top row of Figure 6. These are comparable quantities since $\mathrm{CFI}_{\mathrm{mult}}^{j}$ is equal to the $j$th coefficient in a log-contrast model when it is correctly specified and no zeros are present. We omit plotting the $\ell^{1}$-penalized log-contrast estimates as these are broadly similar to the standard log-contrast estimates in this example. We also plot a marginal effect of $L=\log(Z^{j}/(1-Z^{j}))$ which corresponds to predicting $Y$ using OLS on $L$ and correcting for the presence of zeros as we mention in Remark 1. The results with corresponding $95\%$-confidence intervals are given in the top row of Figure 6; top $10$ most significant species sorted according to significance of $\mathrm{CFI}_{\mathrm{mult}}$ (left) and according to log-contrast coefficients (right).
Figure 6: Effect estimates for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{mult}}$ (top left) and a log-contrast model with pseudocount equal to minimum non-zero observation of $Z$ over $2$ (top right). The labels are colored by the size of the $p$-value. The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model. The bottom left plot shows the $p$-values from different methods for the three most significant species from each method (text color only changes to enhance readability). We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros. The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the top right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments. We see that the relative influences of the different species is highly contingent on the pseudocount.
@@
While the effects overlap for some of the features, there are also significant differences. To further illustrate these differences, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the bottom left of Figure 6. From these $p$-values, we can see that the log-contrast and $\mathrm{CKE}$ are highly correlated and appear to be selecting similar species. This is despite the fact that the log-contrast model is supposed to model the continuous effect of each species. The similarities to the $\mathrm{CKE}$ hence indicates (likely due to the pseudocounts) that the log-contrast coefficients are instead capturing the effects of presence and absence of species. To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the bottom right of Figure 6. This further highlights the dependence of the log-contrast coefficients on the used pseudocount. The fact that the perturbation-based effects explicitly account for zeros and allow us to distinguish between continuous effects (e.g., $\mathrm{CFI}_{\mathrm{unit}}$ or $\mathrm{CFI}_{\mathrm{mult}}$) and presence/absence effects $\mathrm{CKE}$ are an important feature of the perturbation-based approach. In Section S4.7 of the supplementary material we provide an additional plot like the top row of Figure 6 which illustrates the significant hits of $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
5 Theoretical analysis of estimators
In this section we prove asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (16). These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere (e.g., Klyne and Shah, 2023; Chernozhukov et al., 2018). We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.
We start from a given perturbation $\psi$ and a corresponding reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ which reparametrizes $Z$ into $(L,W):=(\phi^{L}(Z),\phi^{W}(Z))$ and assume that $(L,W)$ satisfy (2) if $\psi$ is a binary perturbation and that $\phi$ is derivative-isolating if $\psi$ is a directional perturbation. Our goal is now to use the estimation procedures discussed in Section 3 to estimate the average perturbation effects (i.e., $\tau_{\psi}$ or $\lambda_{\psi}$) consistently and construct asymptotically valid confidence intervals for the estimates. To construct asymptotically valid confidence intervals, we estimate both the target parameter and its variance and then show that the resulting standardized estimators are asymptotically normal with a uniform convergence rate over the class of possible distributions of the data. Distributionally uniform results are desirable here as they ensure uniform coverage guarantees of the confidence intervals. The following result shows uniform asymptotic normality of each of the three estimation procedures.
Theorem 1 (Uniform asymptotic normality).
  1. (a)
    Nonparametric directional effect estimation: Assumption 1 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\tau},\widehat{\sigma}^{2}_{\tau})$ from Algorithm 1.
  2. (b)
    Nonparametric binary effect estimation: Assumption 2 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\lambda},\widehat{\sigma}^{2}_{\lambda})$ from Algorithm 2.
  3. (c)
    Effect estimation using a partially linear model: Assumption 3 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\theta},\widehat{\sigma}^{2}_{\theta})$ from Algorithm 3.
The result is proven in Section S5 of the supplementary material. In the following three sections, we go over each of the settings (a), (b) and (c) individually and provide the required assumptions and detailed algorithms for the target parameter and variance. For the remainder of this section, we omit $\psi$-subscripts for simplicity and use the following notation to help us express the uniformly asymptotic results rigorously. Let $(X_{P,n})_{n\in\mathbb{N},P\in\mathcal{P}}$ denote a family of sequences of random variables indexed by $\mathcal{P}$. For a nonnegative sequence $(a_{n})_{n\in\mathbb{N}}$ we write $X_{P,n}=o_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ it holds that
5.1 Nonparametric estimation for directional effects
We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem 1 (a). A high-level overview of this estimation was given in Section 3.2.2. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S3.4 in the supplementary material). We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm 1 using a user-specified number of splits $K$ in the cross-validation.
Algorithm 1 Nonparametric estimation of $\tau$
4:for $k=1,\dots,K$ do
6:     Fit nuisance function estimates $\widehat{\eta}_{k}$ using $I_{-k}$.
7:     for $r=0,1$ do
To obtain the result in Theorem 1 (a), we require that the estimation methods $\widehat{f}$ and $\widehat{\rho}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption 1.
  1. (a)
  2. (b)
  3. (c)
  4. (d)
  5. (e)
  6. (f)
5.2 Nonparametric estimation for binary effects
We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem 1 (b). A high-level overview of this estimation was provided in Section 3.2.1. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$. We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm 2 using a user-specified number of splits $K$ in the cross-validation.
Algorithm 2 Nonparametric estimation of $\lambda$
3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
9:Set $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}1-L_{i}$.
10:Set $\widehat{\kappa}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}$.
To obtain the result in Theorem 1 (b), we require that the estimation methods $\widehat{f}$ and $\widehat{\pi}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption 2.
  1. (a)
  2. (b)
  3. (c)
  4. (d)
  5. (e)
5.3 Partially linear model for binary/directional effects
Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem 1 (c). A high-level overview of this estimation was provided in Section 3.2.3. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively. We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm 3 using a user-specified number of splits $K$ in the cross-validation.
Algorithm 3 Partially linear estimation of $\theta$ (where $\theta=\lambda$ or $\theta=\tau$)
3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
9:Set $\widehat{J}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{J}_{k}$
11:for $k=1,\dots,K$ do
To obtain the result in Theorem 1 (c), we require that the estimation methods $\widehat{m}$ and $\widehat{g}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption 3.
$\mathbb{E}_{P}[Y\,|\,L,W]=\theta_{P}L+h_{P}(W).$
  1. (a)
  2. (b)
  3. (c)
  4. (d)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{4}=o_{\mathcal{P}}(1)$
  5. (e)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{g}(W_{i})-g_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
  6. (f)

### v4

@@ 4.1.1 Effects of diversity on income in semi-synthetic US census data
We now use this raw individual-level data to construct community-level datasets by artificially grouping individuals together. The grouping is deliberately constructed in a way such that diversity affects income and this relationship is confounded by the remaining variation of the racial composition. This semi-synthetic example is only intended as an illustration. A genuine analysis of this data would require knowledge of the true community structure, which is not available in the data extract we are using here. We emphasize that even if the individual-level data is available, one may still be interested in effects of the racial composition of the community an individual lives in on that individual’s income. For such an analysis, one can add the racial composition of an individual’s community as a variable and then perform the same analysis as here but now on the individual-level data while adjusting for individual-level confounding factors. Importantly, even for such an individual-level analysis, the simplex confounding can bias the results in the same way as we illustrate here. We provide an example of this bias occurring on the individual-level effect in Section S5.5.
@@
We group all individual-level observations into groups of around $50$ individuals. We then create a new response variable $Y\in[0,1]$ by averaging the individual-level compensation in each group. Similarly, we create a predictor variable $X\in\mathbb{R}\times[0,1]\times\{1,\dots,16\}$ by averaging age and sex and selecting the majority class of the categorical variable education (breaking ties randomly). Finally, we create a compositional predictor $Z\in\Delta^{2}$ by computing the proportions of ‘White’, ‘Black’ and ‘Other’ individuals in each group.
@@
To ensure, a (partially) known effect of increased racial diversity on compensation, we construct three categories of groups (0, 1 and 2). We construct these categories based on either education or compensation as follows: Groups in category 1 are sampled to be similarly diverse and with lower educational (compensation) levels than the total population and category 2 groups are sampled to be more diverse and with higher educational (compensation) levels than the total population. The groups in category 0 consist of groups constructed randomly from the remaining individuals.
@@
We then sample groups from each category, using a weighted sampling approach described in detail in Section S5.3. This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation. We expect that – by construction – there is a positive effect of increased diversity on compensation because education is positively associated with compensation. We visualize the distribution of the data and this confounding in Section S5.4.
@@
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use one minus the Gini coefficient $L=1-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 2, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in Section 3.2.3) are provided in Table 2 where random forests are used for nuisance function estimation when including additional covariates. The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded e.g., educational levels may well be correlated with race and average compensation in actual communities.
@@
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.074$ $(-0.046,0.194)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
@@ 4.1.2 Effects of diversity on grades in New York schools
In this section we analyze data from a 2018 Kaggle competition containing data on schools in New York (PASSNYC, 2018). The data contains average grades in English Language Arts (ELA) and in math for 1217 schools along with additional background information on the school such as a summary of the economic need of the students, percentage of students with English as second language, racial composition of the school and various indicators of school performance (for a full list of control variables and our pre-processing of these, see Section S5.6). We let $X$ denote all background variables except racial composition which we denote by $Z$. Race is given as whole percentages of ‘Asian’, ‘Black’, ‘Hispanic’ and ‘White’ students which we rescale so that $Z\in\Delta^{3}$. We are interested in the effect of diversity, again measured by one minus the Gini coefficient, on the average grade of the schools (either ELA or math). We repeat the analysis in Section 4.1.1 on this real data and the results can be seen in Table 3.
@@
$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.348}}$ $(0.108,0.589)$ $0.101$ $(-0.171,0.373)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.161}}$ $(0.002,0.320)$ $0.042$ $(-0.171,0.255)$
@@ 4.2 Variable influence measures when predicting BMI from gut microbiome
We now analyze different types of average perturbation effects and compare them with the commonly used log-contrast based analysis. We fit both a standard log-contrast regression and an $\ell^{1}$-penalized log-contrast regression and then use the coefficients as feature influences. Since the log-contrast model is only defined on the open simplex, that is, it does not permit zeros in the data, one usually adds a ‘pseudocount’ to each observation before rescaling to the simplex. We follow a standard recommendation and use the minimum non-zero observation of $Z$ divided by $2$ (this is not required for the perturbation-based measures). To illustrate the difference between the approaches, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the left of Figure 3. From these $p$-values, we can see that the log-contrast and $\mathrm{CKE}$ are highly correlated and appear to be selecting similar species. This is despite the fact that the log-contrast model is supposed to model the continuous effect of each species. The similarities to the $\mathrm{CKE}$ hence indicates (likely due to the pseudocounts) that the log-contrast coefficients are instead capturing the effects of presence and absence of species. To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the right of Figure 3. This further highlights the dependence of the log-contrast coefficients on the used pseudocount. The fact that the perturbation-based effects explicitly account for zeros and allow us to distinguish between continuous effects (e.g., $\mathrm{CFI}_{\mathrm{unit}}$ or $\mathrm{CFI}_{\mathrm{mult}}$) and presence/absence effects $\mathrm{CKE}$ are an important feature of the perturbation-based approach. In Section S5.7 we provide additional plots which compare the significant hits of $\mathrm{CFI}_{\mathrm{mult}}$, log-contrast estimates and a naive marginal analysis and illustrate the differences between $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
Figure 3: The left plot shows the $p$-values from different methods for the three most significant species from each method (text color only changes to enhance readability). We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros. The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments. We see that the relative influences of the different species is highly contingent on the pseudocount.